OpenStudy (anonymous):

please help @zzrock3r

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

Dully noted, there is no question here.

OpenStudy (anonymous):

hmmm -.- what did you do to the question! put your pencils up!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Okay so that is password protected.

OpenStudy (anonymous):

whats the password?!

OpenStudy (anonymous):

help us. help you!

OpenStudy (anonymous):

i need helo on those questions 2,3,4 and the password is osho

OpenStudy (anonymous):

ow! my head HuRtS! i dont know.. sorry

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

I am not sure how to open that, nor do I really want to. But just ask the question here.

OpenStudy (anonymous):

what have i done wrong ? i am so sorry

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

i know you are angry

OpenStudy (anonymous):

please just forgive .

OpenStudy (zzr0ck3r):

lol why would I be angry?

OpenStudy (zzr0ck3r):

I just said to ask the question :)

OpenStudy (anonymous):

the question is in the file ,

OpenStudy (anonymous):

and the password is osho

OpenStudy (zzr0ck3r):

I don't know how to open the file, and I don't know what the file is, and I don't like opening things when I don't know where they come from. If it is a math question, there is no reason you can't post it here?

OpenStudy (zzr0ck3r):

It does not ask for a password when I try and open it, it says its corrupt or protected, but there is no place for a password.

OpenStudy (anonymous):

(a) Let X ε Rn. Show that the set B(X, ε ) is open. (b) Let X be a complete metric space and {On} is countable collection of dense open subset of X. Show that On is not empty.

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

What is your definition of open?

OpenStudy (anonymous):

2)Let (X, d) and (Y, d) be metric spaces and f a mapping of X into Y. Let τ1 and τ2 be the topologies determined by d and d1 respectively. Then f(X, τ) (y, τ) is continuous if and only if ; that is if x1, x2, . . . , xn, . . . , is a sequence of points in (X, d) converging to x, show that the sequence of points f(x1), f(x2), . . . , f(xn), . . . in (Y, d) converges to x.

OpenStudy (anonymous):

an open interval is an open set. interval without its boundary points

OpenStudy (zzr0ck3r):

How do you define a boundary point?

OpenStudy (zzr0ck3r):

we must show that your ball does not contain any boundary points

OpenStudy (zzr0ck3r):

What would it mean if it did?

OpenStudy (anonymous):

boundary point to me is the being or the end of a thing . in inequality, it is denoted by < or =, and > or = , or [ ]

OpenStudy (zzr0ck3r):

well you are kind of defining what we are asked to prove.

OpenStudy (zzr0ck3r):

A point \(x\) is a boundary point for set \(A\) if \(\forall\) \(\delta>0\) we have \(B_{\delta}(x)\cap A\ne \emptyset\) and \(B_{\delta}(x)\cap A^C\ne \emptyset\)

OpenStudy (zzr0ck3r):

Does this make sense? This says for any neighborhood of \(x\) we have a point in A and a point out of A.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but should it be out of a?

OpenStudy (zzr0ck3r):

Well what you want to show is that \(X\) contains non of its boundary points. So suppose it does. What does that mean?

OpenStudy (anonymous):

it means it is close

OpenStudy (anonymous):

right?

OpenStudy (zzr0ck3r):

brb I need to take my wife to a party. I might be half hour or so.

OpenStudy (zzr0ck3r):

Its only closed if it contains all of them. Suppose it contains one, then by the def I gave you, what does that mean?

OpenStudy (anonymous):

it means it is have open and half close

OpenStudy (zzr0ck3r):

what is?

OpenStudy (zzr0ck3r):

I am here now.

OpenStudy (zzr0ck3r):

Suppose to the contrary that \(B(x, \epsilon)\) contains one of its boundary points, lets call it \(a\). Then \(a\in B(x,\epsilon)\). Now consider \(\delta=\min(d(a, x), d(a, \epsilon))\). Then \(B(a, \delta)\subseteq B(x, \epsilon)\) contradicting the assumption that \(x\) was a boundary point.

OpenStudy (zzr0ck3r):

Thus \(B(x, \epsilon)\) is open because it does NOT contain any of it's boundary points.

OpenStudy (zzr0ck3r):

What does it mean for a metric to generate a topology?

OpenStudy (zzr0ck3r):

For you to understand the second question. You need to know what a metric space is and what does it mean to generate a topology from a metric, in order to understand that you will need to understand what a topology is, and what a basis and sub basis are for a topological space and what it means to be an arbitrary union, in order to understand that you will need to understand what a ray is. Then you will need to know what it means for a sequence of points in some product space to converge, and you must understand the metric definition of continuity, which means you must understand a metric. Do you understand all of these things?

OpenStudy (anonymous):

i know a bit of them but not that good at it

OpenStudy (zzr0ck3r):

Feel free to ask me anything

OpenStudy (anonymous):

5(a) Prove that for any y, z ε , max(y, z) = ½[y + z+ |y-z|], min(y, z) = ½[y + z- |y-z|]. can you try to prove that for me ?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

any \(y,z\in?\)

OpenStudy (anonymous):

of R

OpenStudy (zzr0ck3r):

Suppose w.l.o.g. that \(x<y\). Then \[\max(x,y) = y = \dfrac{x}{2}+\dfrac{y}{2}-\dfrac{x}{2}+\dfrac{y}{2}=\dfrac{1}{2}[x+y-x+y]=\dfrac{1}{2}[x+y+|y-x|]\] Similarly for \(\min\)

OpenStudy (zzr0ck3r):

now you write up a similar proof for the min case, and show it here.

OpenStudy (anonymous):

ok but please just help me so that i don't do any errors

OpenStudy (anonymous):

but sir, why did you use x,y. thought the given function is y,z e R

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

The fact that you don't know that it does not matter if I call it \(z\) or \(x\) really scares me and again begs the question. What are you doing?

OpenStudy (anonymous):

i am trying to know what you did

OpenStudy (michele_laino):

please see the complete solution @GIL.ojei

OpenStudy (michele_laino):

by definition of absolute value, we can write this: \[\Large \begin{gathered} y \geqslant z, \Rightarrow \left| {y - z} \right| = y - z \hfill \\ \hfill \\ y \leqslant z \Rightarrow \left| {y - z} \right| = - \left( {y - z} \right) = - y + z \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

furthermore, those subsequent expressions, namely "max" and "min" are defined for all elements of R, so there is no need for further restrictions.

OpenStudy (michele_laino):

as I said before, we have to distinguish two cases, since R is a totally ordered set. First case: \[y \geqslant z\]

OpenStudy (michele_laino):

then \[y - z \geqslant 0\], so by definition of absolute value, we get: \[\left| {y - z} \right| = y - z\]

OpenStudy (anonymous):

yes, i know that is true

OpenStudy (michele_laino):

then I can replace \[\left| {y - z} \right|\] with \[y - z\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

so I can write this: \[\max \left( {y,z} \right) = \frac{1}{2}\left\{ {y + z + \left| {y - z} \right|} \right\} = \frac{1}{2}\left\{ {y + z + y - z} \right\} = \frac{{2y}}{2} = y\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

furthermore I can write this: \[\begin{gathered} \min \left( {y,z} \right) = \frac{1}{2}\left\{ {y + z - \left| {y - z} \right|} \right\} = \frac{1}{2}\left\{ {y + z - \left( {y - z} \right)} \right\} = \hfill \\ \hfill \\ = \frac{1}{2}\left\{ {y + z - y + z} \right\} = \frac{{2z}}{2} = z \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now we have to consider the other case, namely: \[y \leqslant z\]

OpenStudy (michele_laino):

by definition of absolute value, we have: \[y - z \leqslant 0 \Rightarrow \left| {y - z} \right| = - \left( {y - z} \right) = - y + z\]

OpenStudy (anonymous):

ok. but i thought you have applied that already

OpenStudy (michele_laino):

no, since in this case is y-z less or equal to zero, whereas in the first case is y-z greater or equal to zero

OpenStudy (michele_laino):

in other words, in this second case, we can replace: \[\left| {y - z} \right|\] with: \[ - y + z\]

OpenStudy (anonymous):

are you saying that we ought to have four proves?

OpenStudy (michele_laino):

and then we can write this: \[\begin{gathered} \max \left( {y,z} \right) = \frac{1}{2}\left\{ {y + z + \left| {y - z} \right|} \right\} = \frac{1}{2}\left\{ {y + z + \left( { - y + z} \right)} \right\} = \hfill \\ \hfill \\ = \frac{1}{2}\left\{ {y + z - y + z} \right\} = \frac{{2z}}{2} = z \hfill \\ \hfill \\ \min \left( {y,z} \right) = \frac{1}{2}\left\{ {y + z - \left| {y - z} \right|} \right\} = \frac{1}{2}\left\{ {y + z - \left( { - y + z} \right)} \right\} = \hfill \\ \hfill \\ = \frac{1}{2}\left\{ {y + z + y - z} \right\} = \frac{{2y}}{2} = y \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

no, I say that we have to prove your statement for two different cases for value of y-z

OpenStudy (michele_laino):

that's all!

OpenStudy (anonymous):

which is the main prove ?

OpenStudy (anonymous):

because all you stated are corect

OpenStudy (anonymous):

ok. now i understand sir

OpenStudy (michele_laino):

:)